The following is a news item reported by Reuters:

WASHINGTON, Jan 29 (Reuters)— Orville Medical Group, a maker of reconstructive implants for knees and hips, on Tuesday filed to sell 4 million shares of common stock.

In a filing with the U.S. Securities and Exchange Commission, it said it plans to use the proceeds from the offering for general corporate purposes, working capital, research and development, and acquisitions.

After the sale there will be about 33.0 million shares outstanding in the Arlington, Tennessee-based company, according to the SEC filing.

Orville shares closed at $17.30 on Nasdaq.

The common stock of Orville Medical Group has a par of $0.04 per share.

Required:

Prepare the journal entry to record the sale of the shares assuming the price existing when the announcement was made and ignoring share issue costs.

Note: If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in whole dollars and not in millions.
Answer #1

step 1) 17.30*4= 69.2

0.04*4= 0.16

69.2-0.16= 69.04


answer

dr. cash                                   69,200,000

   Cr. PIC excess of par                                69,040,000

    Cr. common stock                                          160,000

answered by: anonymous